OpenStudy (toxicsugar22):

can u help me with this one sqrt(x+2)=sqrt(x)+1

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

i could have sworn mininaya write a very nice answer to this one. square both sides

OpenStudy (anonymous):

so then x+2 = x +2sqrt(x)+1 no?

OpenStudy (anonymous):

\[\sqrt{x+2}=\sqrt{x}-1\] \[x+2=(\sqrt{x}+1)^2=\sqrt{x}\sqrt{x} + 2\sqrt{x}+1=x+2\sqrt{x}+1\]\]

OpenStudy (anonymous):

so sqrt(x) = 1/2?

OpenStudy (anonymous):

x = 1/4

OpenStudy (anonymous):

so subtracting x from both sides gives \[2=2\sqrt{x}+1\] subtract 1 gives \[1=2\sqrt{x}\] divide by 2 to get \[\frac{1}{2}=\sqrt{x}\] square again to get \[\frac{1}{4}=x\]

OpenStudy (anonymous):

what light said

OpenStudy (toxicsugar22):

ok thanks